# Functional Liftings of Restricted Geometric Inequalities

### v1

@@ 3. Dimensional Functional Brunn-Minkowksi inequalities in measure spaces
Again, the methods of Theorem 1.1 can be applied to $\mathcal{A}_{i}$ the collection of star bodies and applying Theorem C. ∎
Finally we give a functional version of the recent work of Cordero-Erausquin and Eskenazis, [19, Corollary 2].
Theorem D (Cordero-Erausquin & Eskenazis, [19]).
Let $w\colon[0,\infty)\to(-\infty,\infty]$ be an increasing function such that $t\mapsto w(e^{t})$ is a convex function, and suppose that $\mu$ is the measure on ${\mathbb{R}}^{n}$ given by $d\mu(x)=e^{-w(|x|)}dx$. Suppose additionally that $C\subset{\mathbb{R}}^{n}$ is an origin symmetric convex set and $\varphi\colon C\to[0,\infty)$ is even and concave. Given $\beta>0$, the measure defined by $d\nu_{\beta}(x)=\varphi(x)^{\beta}d\mu(x)$, which is supported in $C$, verifies the following inequality: for any $t\in(0,1)$ and and pair of origin symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$, it holds that
$\nu_{\beta}((1-t)K+tL)\geq\mathcal{M}_{\frac{1}{\beta+n}}^{(t)}(\nu_{\beta}(K),\nu_{\beta}(L)).$
Thanks to Theorem 2.3 and Theorem 2.4, we now show that Theorem D above is equivalent to a functional version involving even unimodal functions (note that the direction from functions to sets is obvious).
Theorem 3.3.
Let $\beta>0$, $\alpha\geq-\frac{1}{\beta+n}$, and functions $h,f,g:\mathbb{R}^{n}\to[0,\infty)$, such that $f,g$ are unimodal and even, and satisfy
$h((1-t)x+ty)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
for $x,y\in\mathbb{R}^{n}$. Then, for $\nu_{\beta}$ satisfying the assumptions of Theorem D,
$\nu_{\beta}(h)\geq\mathcal{M}_{r}^{(t)}\left(\nu_{\beta}(f),\nu_{\beta}(g)\right),$
for $r=\frac{\alpha(\beta+n)}{\alpha+n+\beta}$.
Proof.
Proving the result $\alpha=-\frac{1}{\beta+n}$ and $\mathcal{A}=(\mathcal{A}_{1},\mathcal{A}_{2})$ for $\mathcal{A}_{i}$ the cone of all symmetric convex bodies, and $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ to be $\Phi(x,y)=(1-t)x+ty$ and $C(z)=\{(x,y):\Phi(x,y)=z\}$, and applying Theorem 2.3 and then extending to $\alpha>-\frac{1}{\beta+n}$ by Hölder’s inequality (9)as in the Proof of Theorem 1.1, the result follows.
4. Applications to Barthe’s Reverse Brascamp-Lieb Inequality and Schneider’s Conjecture
Here we focus on applications of our results to inequalities concerning the Lebesgue measure on ${\mathbb{R}}^{n}$. In the first subsection, we will treat Barthe’s reverse Brascamp-Lieb inequality proven in [6, 7], while in the second subsection we will show how our results can be used to understand an older conjecture of Schneider from [65] concerning generalized Brunn-Minkowski-type inequalities.
4.1. Barthe’s Reverse Brascamp-Lieb inequality
The celebrated Brascamp–Lieb inequality from [15], asserts that for $m,n\in\mathbb{N}$, $n_{i}\in\mathbb{N}$ and $c_{i}>0$ such that $c_{1}n_{1}+\ldots+c_{m}n_{m}=n$, $B_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n_{i}}$ surjective linear maps, and $f_{i}\in L^{1}(\mathbb{R}^{n_{i}})$ non-negative, the following inequality is valid
$\int_{\mathbb{R}^{n}}\prod_{i=1}^{m}f_{i}(B_{i}x)^{c_{i}}dx\leq D\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}f_{i}\right)^{c_{i}},$
where the constant can be computed by considering just centered Gaussians (see [40] and also [8] for the characterization of $D<+\infty$). In [6, 7], Barthe proved the following reverse form of Brascamp–Lieb inequality.
Theorem E (Barthe).
Let $m,n$ be non-negative integers, $(c_{i})_{1}^{m}$ be a connections of positive real numbers and $(n_{i})_{1}^{m}$ be a collection of non-negative integers smaller than $n$ such that
$\sum_{i=1}^{m}c_{i}n_{i}=n.$
For $i=1,\dots,m$ let $B_{i}$ be a linear surjective map from $\mathbb{R}^{n}$ onto $\mathbb{R}^{n_{i}}$ and set $f=(f_{1},\dots,f_{m})$, with $f_{i}\colon{\mathbb{R}}^{n_{i}}\to[0,\infty)$. If
$\bigcap_{i=1}^{m}\text{ker}B_{i}=\{0\}.$
then the functional
$I(f)=\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}f_{i}^{c_{i}}(x_{i})\colon\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz.$
satisfies the inequality
(10) $I(f)\geq E\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}f_{i}\right)^{c_{i}},$
where
(11) $\begin{split}E&=\inf\left\{\frac{I(g)}{\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}g_{i}\right)^{c_{i}}}\colon g_{i}\hbox{ centered Gaussian on }\mathbb{R}^{n_{i}},i=1,\dots,m\right\}\\ &=\sqrt{\inf\left\{\frac{\det\left(\sum_{i=1}^{m}c_{i}B_{i}^{*}M_{i}B_{i}\right)}{\prod_{i=1}^{m}(\det(M_{i}))^{c_{i}}}\colon M_{i}\in S^{+}(\mathbb{R}^{n_{i}}),i=1,\dots,m\right\}},\end{split}$
and $S^{+}({\mathbb{R}}^{n_{i}})$ denotes the collection of symmetric positive definite $n_{i}\times n_{i}$ matrices.
Set $c\coloneqq\sum_{i=1}^{n}c_{i}$, and for each $i=1,\dots,m,$, let $t_{i}=\frac{c_{i}}{c}$ and $\tilde{f}_{i}=f_{i}^{c}$. Then we can rewrite the inequality (10) in the form
$\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}\tilde{f}_{i}^{t_{i}}(x_{i})\colon\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz\geq E\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}\tilde{f}_{i}^{\frac{1}{c}}\right)^{ct_{i}}.$
Next, by writing $f$ in place of $\tilde{f}$ above, we have the following equivalent formulation of Barthe’s reverse Brascamp-Lieb inequality (10):
$\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}{f}_{i}^{t_{i}}(x_{i});\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz\geq E\prod_{i=1}^{m}\left(\|{f}_{i}\|_{\frac{1}{c}}\right)^{t_{i}}.$
Theorem 1.4 will follow by combining the next results.
Proposition 4.1.
With the definitions of this section, the map
$\square f(z)\coloneqq\sup_{x\in C(z)}\prod_{i=1}^{n}f_{i}^{t_{i}}(x_{i}),$
where
$C(z)\coloneqq\left\{x\in\mathbb{R}^{n_{1}}\times\cdots\times{\mathbb{R}}^{n_{m}}:\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\right\},$
defines a sup-convolution.
Proof.
The goal is the verify the condition of the Definition 2.2. Notice that monotonicity is obvious, while the superadditivity condition follows from the fact that the geometric mean satisfies a reverse Minkowski inequality. Finally, measurability follows from the fact that $\Phi(x)=\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}$ is a continuous function. ∎
Corollary 4.2.
The reverse Brascamp–Lieb inequality (10) is equivalent to the following geometric inequality:
(12) $\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\ \mathcal{M}_{0}^{(c)}\left(|A_{i}|_{n_{i}}\right).$
for sets $A_{i}\subset{\mathbb{R}}^{n_{i}}$.
Proof.
We start with the forward direction. By taking $f_{i}=\mathbbm{1}_{A_{i}}$, we have $\sup_{x\in C(z)}\mathcal{M}_{0}^{(c)}(f_{i}(x_{i}))=\mathbbm{1}_{\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}}$, whereby inequality (10) implies
$\int_{{\mathbb{R}}^{n}}\sup_{x\in C(z)}\mathcal{M}_{0}^{(c)}(f_{i}(x_{i}))dz=\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\prod_{i=1}^{m}\int_{{\mathbb{R}}^{n_{i}{}}}f_{i}=E\ \mathcal{M}_{0}^{(c)}(|A_{i}|_{n_{i}}).$
Conversely, consider the sup-convolution of Proposition 4.1 with $t_{i}\coloneqq\frac{c_{i}}{\|c\|_{1}}$ for $\|c\|_{1}\coloneqq\sum_{i=1}^{m}c_{i}$ and note that by Theorem 2.4 and the homogeneity of the inequality, in order to prove
(13) $\int_{{\mathbb{R}}^{n}}\square f\geq E\ \mathcal{M}_{0}^{(t)}\left(\|f_{i}\|_{{\|c\|_{1}}^{-1}}\right),$
it suffices to check the inequality for $f_{i}=\mathbbm{1}_{A_{i}}$, in which case the inequality is just (12). Now apply (13) to the function $f_{i}=f_{i}^{\|c\|_{1}}$, to obtain
$\int_{{\mathbb{R}}^{n}}\sup_{x\in C(z)}\prod_{i=1}^{m}f^{c_{i}}_{i}(x_{i})dz\geq E\ \prod_{i=1}^{m}\left(\int_{{\mathbb{R}}^{n_{i}}}f_{i}^{\frac{\|c\|_{1}}{\|c\|_{1}}}\right)^{t_{i}\|c\|_{1}}=E\ \prod_{i=1}^{m}\left(\int_{{\mathbb{R}}^{n_{i}}}f_{i}\right)^{c_{i}},$
and we recover (10), as desired. ∎
Remark 4.3.
Such an approach characterizes the finiteness of the constant (11) associated to the reverse Brascamp-Lieb, when one wants to restrict to some class of functions $\mathcal{F}(\mathcal{A})$, to be equivalent with the finiteness of the geometric constant in the inequality (12) for the restricted class $\mathcal{A}$, as it can be seen by repeating verbatim the above proof for this class.
4.2. A Functional Version of Schneider’s Conjecture
In [65] Schneider considered the following generalization of Minkowski difference (called the $m$-order difference body): given a convex body $K\subset\mathbb{R}^{n}$, and $n\geq 2,m\geq 1$, define the $m$-th order difference body of $K$ via
$D^{m}(K)=\{x=(x_{1},\dots,x_{m})\in(\mathbb{R}^{n})^{m}\colon K\cap(x_{1}+K)\cdots\cap(x_{m}+K)\neq\emptyset\}.$
In [65], Schneider established the following variant of the celebrated Rogers-Shephard inequality:
(14) $|D^{m}(K)|_{nm}|K|^{-m}\leq\binom{n(m+1)}{n}$,
with equality only for $K$ a simplex in ${\mathbb{R}}^{n}$. Schneider additionally proved the following generalization of the Brunn-Minkowski inequality for $n=2$ and $m$ arbitrary. For any convex body $K\subset{\mathbb{R}}^{2}$, it holds that
(15) $|D^{m}(K)|_{2m}|K|_{2}^{-m}\geq(m+1)^{2},$
with equality if and only if $K$ is centrally symmetric. Note that for $n\geq 1$ and $m=1$, the above reduces to the Brunn-Minkowski inequality applied to KKK-K.
Various extensions of known geometric operators were inspired by the construction appearing in [65]; this theory been rapidly developing (see [60, 31, 35, 37, 34, 36, 32]). In particular a functional version of the inequality (14) was proven in [34], while a measure theoretic variant was established in [60].
The following fascinating question was left open in [65].
Conjecture 4.4 (Schneider’s conjecture (1970)).
For $n\geq 3$ and $m\geq 2$, is $|D^{m}(K)|_{nm}|K|^{-m}$ minimized by ellipsoids? In order words, is it necessarily the case that
$\min_{K\subset{\mathbb{R}}^{n}}|D^{m}(K)|_{nm}|K|^{-m}=|D^{m}(B)|_{nm}|B|^{-m}$
whenever $n\geq 3,m\geq 2$? Here $B\subset{\mathbb{R}}^{n}$ is the $\ell_{2}$ unit ball.
It is worth noting that this question has an intimate connection with a long standing conjecture of Petty appearing in [57], but we do no elaborate on this connection here. Very recently, in [32], Conjecture 4.4 was proven with $D^{m}K$ replaced by $(D^{m}K)^{\circ}$, the polar set, where the maximizers were shown to be ellipsoids. Thus, pairing this result with the bounds for the volume product, Conjecture 4.4 is known to hold up to an absolute constant. Additionally, the authors proved a functional version of this dual inequality using the functional Blaschke-Santalo inequality and Barthe’s reverse Brascamp-Lieb inequality.
Proposition 4.5.
Let $m,n$ be positive integers and $\mathcal{A}$ be the collection of convex bodies in ${\mathbb{R}}^{n}$. Consider $\bar{f}=(f,\dots,f)$, with $f\in\mathcal{F}({\mathcal{A}})$ being non-negative. Then the map $\square_{m}\bar{f}\colon({\mathbb{R}}^{n})^{m}\to[0,\infty)$ defined by
$\square_{m}\bar{f}(z)\coloneqq\left(\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f(x-z_{i})\right)^{\frac{1}{m+1}}$
is a sup-convolution.
Proof.
The monotonicity is clear from the definition of the supremum, which verifies (1) from Definition 2.2. Suppose that $\bar{f}=(f,\dots,f),\bar{g}=(g,\dots,g)$ are step functions for which $f(x_{i})\leq f(y_{i})$ for all $y_{i}\in{\mathbb{R}}^{n}$ whenever $x_{i}\in{\mathbb{R}}^{n}$ is such that $g(x_{i})>0$. We need to prove that $\square_{m}(f+g)\leq\square_{m}f+\square_{m}g$. Should $g\equiv 0$, then the result is obvious. Suppose that $g\not\equiv 0$, and that $z$ is such that $\square g(z)>0$, and choose $\varepsilon>0$ sufficiently small such that $\square g(z)-\varepsilon>0$. Next, select $x\in{\mathbb{R}}^{n}$ such that
$g(x)g(x-z_{1})\cdots g(x-z_{m})\geq\square_{m}\bar{g}(z)-\varepsilon.$
Necessarily, for each $i$, this we obtain $f(x-z_{i})>0$, whence $f(x-z_{i})\geq f(y-z_{i})$ holds for all $y$; hence,
$f(x)f(x-z_{1})\cdots f(x-z_{m})=\square_{m}\bar{f}(z).$
Letting $z_{0}=0$, we obtain
$\square_{m}\bar{f}(z)+\square_{m}\bar{g}(z)$ $\leq g(z)g(x-z_{1})\cdots g(x-z_{m})+f(x)f(x-z_{1})\cdots f(x-z_{m})$
$\leq\prod_{i=0}^{m}\left(g(x-z_{i})+f(x-z_{i})\right)+\varepsilon$
$\leq\sup_{y\in{\mathbb{R}}^{n}}\prod_{i=0}^{m}\left(g(y-z_{i})+f(y-z_{i})\right)+\varepsilon$
$=\square_{m}(f+g)(z)+\varepsilon.$
Definition 2.2 (2) follows from the inequality $\prod_{i=0}^{m}a_{i}+\prod_{i=0}^{m}b_{i}\leq\prod_{i=0}^{m}(a_{i}+b_{i})$ for $a_{i},b_{i}\geq 0$. Taking $\varepsilon\to 0$ completes the proof. Finally, the condition (3) follows from the measurability of $D^{m}(K)$.
Corollary 4.6.
Let $m\geq 2,n\geq 3$. The geometric inequality in Conjecture 4.4 is equivalent to the functional inequality: given $\bar{f}=(f,\dots,f)$, for $f\colon{\mathbb{R}}^{n}\to[0,\infty)$ unimodal, it holds that
(16) $\int_{(\mathbb{R}^{n})^{m}}\square_{m}\bar{f}(z)\geq c(m,n)\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m},$
where $c(m,n)=|D^{m}(B)|_{nm}|B|^{m}$ and $B$ is the $\ell_{2}$ unit ball.
Proof.
To prove that the functional inequality implies Schneider’s geometric conjecture, take $f=\mathbbm{1}_{K}$, so that $\square f(z)=\mathbbm{1}_{D^{m}(K)}$, the integral inequality is then
$|D^{m}(K)|_{nm}\geq c(m,n)|K|^{m},$
which is exactly Schneider’s conjectured inequality. If Schneider’s conjectured inequality holds, then the inequality holds for $f$ the indicator of a convex set, noting that the inequality is homogeneous we have the result for $f$ a scaled step function. Interpreting $\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m}=\mathcal{M}_{1}^{(1/2)}(\|f\|_{\frac{1}{m}})$ where $\mathcal{M}_{1}^{(1/2)}(u)=u$ as it is the mean of one value, and applying Theorem 2.4 the result follows. ∎
We mention in passing that the following operation considered in [34] can be confirmed by analogous methods to be a sup-convolution as well.
Definition 4.7.
For $\mathcal{A}$ the space of convex bodies and $f_{i}:\mathbb{R}^{n}\to[0,\infty)$ belonging to $\mathcal{F}(\mathcal{A})$, define $\square{f}:(\mathbb{R}^{n})^{m}\to[0,\infty)$
$\square f(z)\coloneqq\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f_{i}(x-z_{i}),$
where $z_{0}\coloneqq 0$ for notational brevity.
We note that Schneider’s conjecture is confirmed for convex sets $K\subseteq\mathbb{R}^{n}$ when $n=2$ and $m\geq 1$ is arbitrary, hence (16) is a valid functional inequality for $n\leq 2$ by (15), which constitutes a generalization of the Prékopa-Leindler inequality (3) when $n=2$ for the function
$\square_{1}f(z)=\sup_{z=x-y}f(x)f(y).$
Theorem 4.8.
Let $n=2$, $m\geq 1$. Then, for any unimodal function $f\colon{\mathbb{R}}^{n}\to[0,\infty)$, we have
$\int_{(\mathbb{R}^{2})^{m}}\square_{m}f(z)\geq(m+1)^{2}\left(\int_{\mathbb{R}^{2}}f^{\frac{1}{m}}\right)^{m}.$
Moreover, this functional inequality is equivalent to (15).
Proof.
To prove that the functional inequality implies the geometric inequality (15), it is enough choose $f=\mathbbm{1}_{K}$, whereby $\square_{m}f=\mathbbm{1}_{D^{m}(K)}$. Conversely, combining the inequality (15) with Theorem 2.4 completes the proof. ∎
@@ 5. Functional $L_{p}$-Brunn-Minkowski inequalities
We remind the reader that we take a convex body to be a compact convex set with non-empty interior. For convex bodies containing the origin in their interior, $K_{0}$ and $K_{1}$, $t\in(0,1)$ and $p\in[0,\infty)$ one defines the $L_{p}$-Minkowski convex combination of $K_{0}$ with $K_{1}$, denoted $(1-t)\cdot K_{0}+_{p}t\cdot K_{1}$, via the Wülff shape of $p$-th average of the support functions of the $K_{i}$. More explicitly, for a set $A\subseteq\mathbb{R}^{n}$ define $h_{A}:\mathbb{R}^{n}\to\mathbb{R}$ by
$h_{A}(z)\coloneqq\sup_{x\in A}\langle x,z\rangle.$
@@
As proved in [51, Lemma 1-2] this interpretation of the $L_{p}$-Minkowski sum agrees with (17) when $p>1$ and when $A,B$ are convex bodies containing the origin in their interiors. Therein, the authors established the following general form of the $L_{p}$-Brunn-Minkowski inequality for Borel sets for $p>1$:
@@
Since $(1-t)A+tB\subseteq(1-t)A+_{p}tB$ for $p>1$ (a non-trivial inclusion proved in [51, Lemma 3]), inequality (18) follows from the classical Brunn-Minkowski inequality (1) and a homogeneity argument.
@@ 5.1. The case $p\in[0,1)$
Theorem F (Chen, Huang, Li, & Liu, [16]).
There exists a $p_{0}\in(0,1)$, depending on the dimension, such that for all $p\in(p_{0},1)$ and for any pair of origin-symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$ and any $t\in[0,1]$, it holds that
@@
Let $p_{0}$ be as in Theorem F and let $p\in(p_{0},1)$. Let $\alpha\geq-\frac{p}{n}$, and $h,f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be a triple of functions, with $f,g$ being even and unimodal, such that
@@
Theorem H (Livshyts, Marsiglietti, Nayar and Zvavitch, [41]).
@@
holds for all $t\in[0,1]$ and $A,B$ symmetric convex bodies.
@@
The log-Brunn-Minkowski conjecture is true if and only if its functional extension holds. Moreover the log-Brunn-Minkowski inequality implies that the $L_{p}$-Borell-Brascamp-Lieb inequalities hold for all unimodal symmetric functions with respect to even log-concave measures.
@@
then for a symmetric log-concave measure $\mu$,
@@
The next result can be deduced from the articles [63, 64].
Theorem I (Saroglou, [63, 64]).
Let $\mu$ be a $\log$-concave measure with an unconditional density on Then, for any pair of unconditional convex bodies $K,L\subset{\mathbb{R}}^{n}$ and any $t\in[0,1]$,
$\mu(K^{1-t}L^{t})\geq\mathcal{M}_{0}^{(t)}(\mu(K),\mu(B)).$
Combining Theorem 5.4, with Theorem I and Corollary 5.6, we get Theorem 1.3

### v3

@@ 3. Dimensional Functional Brunn-Minkowksi inequalities in measure spaces
The methods of Theorem 1.1 can be applied to $\mathcal{A}_{i}$ the collection of star bodies and applying Theorem C. ∎
When restricted to radially decreasing functions and $\alpha>0$, this becomes [1, Theorem 1.8].
@@ 4. Functional $L_{p}$-Brunn-Minkowski inequalities
We remind the reader that we take a convex body to be a compact convex set with non-empty interior. For convex bodies containing the origin in their interior, $K_{0}$ and $K_{1}$, $t\in(0,1)$ and $p\in[0,\infty)$ one defines the $L_{p}$-Minkowski convex combination of $K_{0}$ with $K_{1}$, denoted $(1-t)\cdot K_{0}+_{p}t\cdot K_{1}$, via the Wülff shape of $p$-th average of the support functions of the $K_{i}$. More explicitly, for a set $A\subseteq\mathbb{R}^{n}$ define (the support function) $h_{A}:\mathbb{R}^{n}\to\mathbb{R}$ by
(19) $h_{A}(z)\coloneqq\sup_{x\in A}\langle x,z\rangle.$
@@
As proved in [43, Lemma 1-2] this interpretation of the $L_{p}$-Minkowski sum agrees with (20) when $p>1$ and when $A,B$ are convex bodies containing the origin in their interiors. Therein, the authors established the following general form of the $L_{p}$-Brunn-Minkowski inequality for Borel sets for $p>1$:
@@
Since $(1-t)A+tB\subseteq(1-t)A+_{p}tB$ for $p>1$ (a non-trivial inclusion proved in [43, Lemma 3]), inequality (21) follows from the classical Brunn-Minkowski inequality (1) and a homogeneity argument.
@@ 4.1. The case $p\in[0,1)$
Theorem D (Chen, Huang, Li, & Liu, [12]).
There exists a $p_{0}\in(0,1)$, depending on the dimension, such that for all $p\in(p_{0},1)$ and for any pair of origin symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$ and any $t\in[0,1]$, it holds that
@@
Let $p_{0}$ be as in Theorem D and let $p\in(p_{0},1)$. Let $\alpha\geq-\frac{p}{n}$, and $h,f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be a triple of functions, with $f,g$ being even and unimodal, such that
@@
Theorem F (Livshyts, Marsiglietti, Nayar and Zvavitch, [33]).
@@
holds for all $t\in[0,1]$ and $A,B$ origin symmetric convex bodies.
@@
The log-Brunn-Minkowski conjecture is true if and only if its functional extension holds. Moreover the log-Brunn-Minkowski inequality implies that the $L_{p}$-Borell-Brascamp-Lieb inequalities hold for all even unimodal functions with respect to even log-concave measures.
@@
then for any even log-concave measure $\mu$,
@@
Remark 4.7.
As we mentioned in the introduction, the log-Brunn-Minkowski conjecture in the unconditional case follows from an application of the Prèkopa-Leindler inequality, in [53]. It is in fact more immediate to see this through the functional analogue we introduced. Indeed, let $f,g:\mathbb{R}^{n}\to[0,+\infty)$ be unconditional unimodal functions and $h:\mathbb{R}^{n}\to[0,+\infty)$ satisfying $h(z)\geq f(u)^{1-t}g(v)^{t}$ for each $z\in\mathbb{R}^{n}$ and $u,v$ such that $z\in\{f\geq f(u)\}^{1-t}\{g\geq g(v)\}^{t}$. Then, if $u,v\in[0,+\infty)^{n}$ and $z:=(u_{1}^{1-t}v_{1}^{t},\ldots,u_{n}^{1-t}v_{n}^{t})$, for each $\theta\in S^{n-1}$ and $\varepsilon_{i}:=sgn(\theta_{i})$, setting $\tilde{u}=(\varepsilon_{1}u_{1},\ldots,\varepsilon_{n}u_{n}),\tilde{v}=(\varepsilon_{1}v_{1},\ldots,\varepsilon_{n}v_{n})$ it holds
$\langle z,\theta\rangle=\sum_{i=1}^{n}\varepsilon_{i}|u_{i}\theta_{i}|^{1-t}|v_{i}\theta_{i}|^{t}\leq\langle\tilde{u},\theta\rangle^{1-t}\langle\tilde{v},\theta\rangle^{t}.$
Since by unconditionality $\tilde{u}\in\{f\geq f(u)\}$ and $\tilde{v}\in\{g\geq g(v)\}$, we conclude that $z\in\{f\geq f(u)\}^{1-t}\{g\geq g(v)\}^{t}$ (equivalently $(u,v)\in C_{0}(z,f,g)$) and thus $h(u_{1}^{1-t}v_{1}^{t},\ldots,u_{n}^{1-t}v_{n}^{t})\geq f(u)^{1-t}g(v)^{t}$. The wanted functional inequality now follows from the multiplicative Prèkopa-Leindler inequality which may be found for example in the paper of Uhrin [57]; this follows by considering $\tilde{f}(t_{1},\ldots,t_{n}):=e^{t_{1}+\ldots+t_{n}}f(e^{t_{1}},\ldots,e^{t_{n}})$, same for $g$ and $h$, and using the classical Prèkopa-Leindler.